OpenStudy (anonymous):

Factor plox x^4-16x^2+24

OpenStudy (popitree):

ax^2+bx+c, try to factorize ac = 1*24 = 24, and with the factors come with 2 numbers when added will give you 16

OpenStudy (popitree):

i mean 2 numbers when multiplied will give u 24 and when added will give u 16

OpenStudy (anonymous):

But the polynomial has a degree of 4 so it can't be that simple

OpenStudy (popitree):

if this not satisfied consider x^2=y, solve for y using Sreedhar Acharya formula, u will get 2 roots, then square root each of them u will get 4 roots

OpenStudy (popitree):

by the way, check if it is 2x^4, otherwise it will be a bit complex root

OpenStudy (anonymous):

Doesn't have a coefficient. This is pretty hard

OpenStudy (anonymous):

It's definitely a complex root. Thanks!

OpenStudy (popitree):

u will get

OpenStudy (popitree):

\[x ^{2} = 8\pm2\sqrt{10}\]